OpenStudy (anonymous):

find the equation of the line containing (-6,5) that is parallel to 6x-2y=-8

OpenStudy (eamier):

the gradient of the new equation should be same as gradient of 6x-2y=-8

OpenStudy (anonymous):

y = mx +b

OpenStudy (eamier):

that is ok. so the m for both equation are same

OpenStudy (eamier):

find the m

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

to find m = -A / B

OpenStudy (anonymous):

m = -6/-2 m=3

OpenStudy (eamier):

ok

OpenStudy (eamier):

can you continue

OpenStudy (anonymous):

line parallel to that is 6x-2y=D , now put x=-6 and y=5 6*-6-2*5=D D=-46 and your line is 6x-2y=-46